# Diffusion-Refined VQA Annotations for Semi-Supervised Gaze Following

### v1

@@ Abstract
Training gaze following models requires a large number of images with gaze target coordinates annotated by human annotators, which is a laborious and inherently ambiguous process. We propose the first semi-supervised method for gaze following by introducing two novel priors to the task. We obtain the first prior using a large pretrained Visual Question Answering (VQA) model, where we compute Grad-CAM heatmaps by ‘prompting’ the VQA model with a gaze following question. These heatmaps can be noisy and not suited for use in training. The need to refine these noisy annotations leads us to incorporate a second prior. We utilize a diffusion model trained on limited human annotations and modify the reverse sampling process to refine the Grad-CAM heatmaps. By tuning the diffusion process we achieve a trade-off between the human annotation prior and the VQA heatmap prior, which retains the useful VQA prior information while exhibiting similar properties to the training data distribution. Our method outperforms simple pseudo-annotation generation baselines on the GazeFollow image dataset. More importantly, our pseudo-annotation strategy, applied to a widely used supervised gaze following model (VAT), reduces the annotation need by 50%. Our method also performs the best on the VideoAttentionTarget dataset.
@@ 1 Introduction
Human gaze behavior is a vital cue for applications such as human-machine interaction [39, 11, 1], social interaction analysis [35, 5], and human intention interpretation [48, 52]. Although eye-tracking glasses can be used to track gaze behavior, they are intrusive, require personalized calibration, and are mostly restricted to laboratory environments. In contrast, the gaze following task [45, 9, 12] predicts a person’s gaze target in an image in-the-wild, by predicting a target heatmap from an input scene image with the person’s head bounding box.
@@
This is the first semi-supervised gaze following method, to the best of our knowledge. Semi-supervised recognition and segmentation methods are typically not directly applicable to gaze following due to task differences. As gaze following models predict a target heatmap instead of classification logits and require an intact scene image as input, methods that operate on the predicted multi-class distribution [57, 26], learn disentangled representations [42, 33], or perform strong data augmentation (e.g.,CutOut[54, 13]) cannot easily be adapted to gaze following. Additionally, for semi-supervised methods that use a teacher model to pseudo-annotate the unlabeled data [30, 59, 29], it is unclear how to design a gaze following teacher model to generate good quality pseudo-annotations.
@@
In summary, we propose the first semi-supervised gaze following method, in which we introduce two novel priors to the task: (1) Prior knowledge from pre-trained vision-language models for initial annotation generation. (2) A novel diffusion-based annotation prior to refine the noisy VL annotations into reliable pseudo-labels.
@@ 2 Prior Work
Gaze Following was first introduced in [45] together with the GazeFollow dataset and a model composed of two separate pathways for encoding gaze orientation and scene saliency information, also adopted by later work [34, 8, 9, 12]. Chong et al. [8] considered out-of-frame targets and also extended the task to videos [9], proposing the VideoAttentionTarget dataset and the VAT model. Later work has improved gaze following models by leveraging monocular depth estimations [12, 25, 60, 21] and human poses [16, 4]. Tu et al. [62] proposed a transformer model for gaze following, while [24, 38] added training losses for numerical coordinate regression and patch-level gaze distribution prediction. Recent works investigate 3D gaze following [22, 20], gaze following for children [58], and object-aware gaze target prediction [61]. All of these methods are fully-supervised.
@@
Semi-supervised Learning. The majority of semi-supervised learning methods adopt a teacher-student framework, which can be categorized into two groups: self-training [30, 7, 15, 54] and consistency regularization [29, 59, 13, 18]. Self-training methods generate pseudo annotations from teacher models trained with labeled data, while consistency regularization applies a consistency loss between the student and teacher model outputs, updating the teacher model gradually during training. As most semi-supervised learning methods are built for recognition or segmentation tasks, they are not directly applicable to gaze following [54, 42, 13, 26] due to task differences. However, some general methods that do not perform task-specific operations [29, 59] can be applied to gaze following with appropriate modifications (details in the Supplementary). In our method, we adopt a self-training pipeline using the VQA prior and diffusion-based annotation prior, and show that we can get further improvements by using applicable consistency regularization methods (e.g. Mean Teacher [59]) to enhance the training of the diffusion model.
@@
Vision-Language Models. Vision-Language (VL) models [63, 27, 51, 31] have succeeded in tasks such as visual grounding, visual question answering (VQA), and image captioning. They also do well in low-shot generalization tasks by “prompting” [43, 37, 31], where appropriate task instructions (usually in the form of text) are given to a pretrained model to generate outputs useful for other tasks. Grad-CAM [50] visualizations on the cross-attention maps localize the linguistic input effect on the visual input [32, 66] thus achieving a degree of interpretability.
@@
Diffusion Models. Denoising diffusion probabilistic models (DDPM) [53, 19] give state-of-the-art results on image synthesis, by capturing intricate data properties [10, 40, 46, 28]. They have also been used in inverse problems [56], image editing [36], and adversarial purification [41]. These works treat images with added noise as intermediate steps of the reverse sampling process, showing that the noise magnitude affects the information retained from input images. In our task, we add appropriate magnitude noise to the Grad-CAM heatmaps to achieve the best trade-off between the VQA prior and annotation prior. Recent work adopts diffusion models in semi-supervised learning. [18] uses a diffusion model as the student model for 3D object detection, whereas [65] trains a diffusion model for image generation with images pseudo-labeled by existing semi-supervised models. In contrast, we use diffusion model as annotation prior, to refine the initial noisy annotations from VQA into high-quality pseudo annotations.
@@ 3.1 Overview
Figure 2: (a) Overall pipeline for semi-supervised training. We compute Grad-CAM heatmaps for a large set of unlabeled images and train the diffusion model with a small human-labeled set. We use the diffusion model to refine the Grad-CAM heatmaps into pseudo-annotations. Both the pseudo-annotations and the human-labeled training set are used to train a gaze following model. (b) Grad-CAM heatmap generation procedure. Given an input image with an overlaid bounding box for the person, we “prompt" a large pretrained VQA model with a gaze question and compute the Grad-CAM heatmap from the answer. (c) Grad-CAM refinement process. We perturb the raw Grad-CAM heatmaps with Gaussian noise and pass through the reverse diffusion process to generate the refined pseudo-annotations.
@@ 3.2 Grad-CAM heatmap extraction
We use OFA [63], an open-source large pretrained VL model as the VQA model. Fig. 2b illustrates the procedure for Grad-CAM heatmap generation. Given an image $\bm{I}$ and the head bounding box of a person $\bm{l}_{h}$, we use Mask R-CNN [17] for person detection and find the bounding box that maximally overlaps with $\bm{l}_{h}$. We overlay this bounding box on top of the input image and give it as input to the VQA model along with a “prompt” in the form of a question: “What is the person in the green bounding box looking at?".
OFA has a transformer-based encoder-decoder structure. The model generates a sequence of words as an answer. We compute a Grad-CAM heatmap $\mathcal{G}=\{\bm{g}_{j}\}_{j=1}^{N_{u}}$ on the decoder cross-attention weights between the input query token, and the image patch tokens based on [6] after selecting the noun from the answer. Details are in the supplementary.
@@ 3.3 Diffusion Model Training
Fig. 3(a) shows the training procedure. At each iteration, we sample a random time step $t$ from $U[1,T]$ to generate a noisy input $\mathbf{h}_{t}$ using Eq. 4. We feed $\mathbf{h}_{t}$ to the diffusion model for single-step denoising, and optimize with the loss of Eq. 5. We condition the diffusion model with the gaze feature $\mathbf{c}$ extracted from the input triplet $\bm{X}$ (input image, head crop, and head location mask) to output a gaze heatmap grounded on the gazing person; the parametrization of the U-Net denoiser becomes $\mathbf{h}_{0}^{\theta}(\mathbf{h}_{t},t,\mathbf{c})$. The feature extractor follows the structure of the VAT model [9], which takes the input triplet $\bm{X}$ and outputs the extracted features concatenated from two pathways for encoding the scene and gaze features. Details of the VAT structure are shown in Supplementary. Following [14], we add the extracted features as conditional features to each layer in the U-Net denoiser with the necessary transposed convolution layers to match the feature size of U-Net. During the reverse sampling process, we used DDIM [55] to sample the next step input using the predicted $\mathbf{h}_{0}$ from the U-Net.
@@ 3.4 Heatmap Refinement using Diffusion Models
The refinement process is shown in Fig. 3(b). Given a noisy Grad-CAM heatmap $\bm{g}$ of an unsupervised input sample $\bm{X}$, we first add an appropriate amount of noise using the forward diffusion process $\hat{\mathbf{h}}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{g}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ according to Eq. 4. The added noise smoothes out the noise artifacts of the Grad-CAM heatmaps, while preserving the information of the highlighted regions. Then, by iteratively applying the steps of the denoising procedure in Eq. 3, we get samples of the final heatmap $\hat{\mathbf{h}}_{0}$ that are geometrically and contextually similar to the human-labeled heatmap $\mathbf{h}_{0}$ while retaining gaze target information from the noisy Grad-CAM outputs.
@@ 4.1 Setting and Implementation Details
Implementation Details. We used a diffusion model with a linear noise schedule and $T{=}500$ steps. The heatmap size was $64{\times}64$. We used a batch size of 80 and a learning rate of $2.5\times 10^{-4}$ in the semi-supervised learning experiments. For training the diffusion model on the pool of labeled data, we used a learning rate of $5\times 10^{-5}$. When refining Grad-CAM heatmaps, we used DDIM [55] for inference, with two inference steps.
@@ 4.2 Semi-supervised Training Results
We used VAT as the student model and compared our method of generating pseudo annotations with several alternatives [^1] Implementation details of the baselines are provided in the supplementary material.: 1) Semi-VAT: the pseudo-annotations for the unlabeled data were generated by the VAT model trained with labeled images. 2) Semi-VAT-GC: We created VAT-GC by adding Grad-CAM heatmaps as conditional inputs to the VAT model and modifying the VAT model accordingly. VAT-GC was trained on the labeled data to generate pseudo annotations for the unlabeled data. 3) VAT-MT: We use the Mean Teacher (MT) method [59] as a representative consistency regularization method to train the VAT model, which updates the teacher model using both labeled and unlabeled data through an exponential moving average (EMA) of student model weights. Here we applied the consistency loss in MT between the heatmap output from the teacher and student models. The first two baselines are self-training methods, while the 3rd belongs to the consistency regularization category.
@@ 4.3.2 Added Noise Magnitude Effects
We further visualized the outputs from the same diffusion model when adding noise at different time steps in Fig. 5. In the top 2 rows, the Grad-CAM heatmaps highlight correct locations, whereas the Grad-CAM heatmaps of the bottom 2 rows are noisy or highlight incorrect locations. In all these cases, when adding noise at the 100th step, the outputs are located on the highlighted region of the Grad-CAM heatmap. In contrast, when adding noise at the 400th step, the large magnitude of added noise leads to outputs similar to sampling directly from Gaussian noise without using the Grad-CAM heatmaps. Adding noise at step 250 appears to be the best trade-off.
@@ 4.3.3 The Effect of the Grad-CAM Heatmaps
In this section, we test how much guidance signal is provided by the VQA priors. This cannot be directly tested in the semi-supervised scenario, so we tested by using the Grad-CAM heatmaps as a direct input to a fully supervised gaze following model. We train the teacher models of Sec. 4.2 using different amounts of labeled data only, and introduce a new baseline: Diffusion-GC, where the diffusion model trained on labeled images was used to refine the Grad-CAM heatmaps computed directly on the test set. Note that VAT-GC and Diffusion-GC require the Grad-CAM heatmaps of the test images during inference, which may not be feasible in an online setting as it necessitates access to a large vision-language model.
Tab. 4 presents the results. VAT-GC outperforms the baseline VAT model in all cases, and Diffusion-GC performs best in the distance metrics. Both the VAT and diffusion model show significant performance boosts when Grad-CAM heatmaps are directly input into the gaze following models, which offers evidence that the VQA Grad-CAM heatmaps bring strong prior knowledge. As illustrated in Sec. 4.2, the lower AUC for Diffusion and Diffusion-GC can be attributed to the AUC on GazeFollow being evaluated with group-level annotations and favoring heatmaps spanning larger areas. We analyzed this in more detail in the supplementary material.

### v2

@@ Abstract
Training gaze following models requires a large number of images with gaze target coordinates annotated by human annotators, which is a laborious and inherently ambiguous process. We propose the first semi-supervised method for gaze following by introducing two novel priors to the task. We obtain the first prior using a large pretrained Visual Question Answering (VQA) model, where we compute Grad-CAM heatmaps by ‘prompting’ the VQA model with a gaze following question. These heatmaps can be noisy and not suited for use in training. The need to refine these noisy annotations leads us to incorporate a second prior. We utilize a diffusion model trained on limited human annotations and modify the reverse sampling process to refine the Grad-CAM heatmaps. By tuning the diffusion process we achieve a trade-off between the human annotation prior and the VQA heatmap prior, which retains the useful VQA prior information while exhibiting similar properties to the training data distribution. Our method outperforms simple pseudo-annotation generation baselines on the GazeFollow image dataset. More importantly, our pseudo-annotation strategy, applied to a widely used supervised gaze following model (VAT), reduces the annotation need by 50%. Our method also performs the best on the VideoAttentionTarget dataset. Code is available at https://github.com/cvlab-stonybrook/GCDR-Gaze.git
Abstract
In this supplementary material, we provide more details about the experiment settings of the main paper (Sec. S6), we describe the baseline implementation details (Sec. S7), we illustrate the details of the Grad-CAM heatmap generation procedure (Sec. S8), we present the results for each video in the video semi-supervised learning task (Sec. S9), we discuss the reason for the lower AUC of the diffusion models in the fully supervised experiments (Sec. S10), we show the result of training with 50% annotations on GazeFollow (Sec. S11) and training with 10% labels on VideoAttentionTarget using the normal train-test split (Sec. S12), we applied our methods to other gaze following models (Sec. S13), we analyzed the effect of the amount of unlabeled data by fixing the amount of labeled data (Sec. S14), we visualize outputs from the diffusion model at intermediate inference steps (Sec. S15), we provide detailed discussions on the available semi-supervised learning methods and their relations to gaze following (Sec. S16) and evaluate the raw Grad-CAM heatmaps when obtained with different extraction methods (Sec. S17).
@@ 1 Introduction
Human gaze behavior is a vital cue for understanding human cognitive processes [95, 96] and for applications such as human-machine interaction [63, 17, 1], social interaction analysis [57, 8], and human intention interpretation [73, 80]. In contrast to eye-tracking glasses which are intrusive and mostly restricted to laboratory environments, the gaze following task [69, 12, 18] predicts a person’s gaze target in an image in-the-wild, by predicting a target heatmap from an input scene image with the person’s head bounding box.
@@
This is the first semi-supervised gaze following method, to the best of our knowledge. Semi-supervised recognition and segmentation methods are typically not directly applicable to gaze following due to task differences. As gaze following models predict a target heatmap and require an intact scene image as input, methods that operate on the predicted multi-class distribution [85, 37], perform strong data augmentation (e.g.,CutOut[82, 19]) cannot easily be adapted to gaze following. Additionally, for semi-supervised methods that use a teacher model to pseudo-annotate the unlabeled data [45, 87, 44], it is unclear how to design a gaze following teacher model to generate good quality pseudo-annotations.
@@
In summary, we propose the first semi-supervised gaze following method, in which we introduce two novel priors to the task: (1) Prior knowledge from pre-trained VQA models for initial annotation generation. (2) A novel diffusion-based annotation prior to refine the noisy VL annotations into reliable pseudo-labels.
@@ 2 Prior Work
Gaze Following was first introduced in [69] together with the GazeFollow dataset and a model composed of two separate pathways for encoding gaze orientation and scene saliency information, also adopted by later work [50, 11, 12, 18]. Chong et al. [11] considered out-of-frame targets and also extended the task to videos [12], proposing the VideoAttentionTarget dataset and the VAT model. Later work has improved gaze following models by leveraging monocular depth estimations [18, 36, 88, 31] and human poses [24, 5]. Tu et al. [91] proposed a transformer model for gaze following, while [35, 61] added training losses for numerical coordinate regression and patch-level gaze distribution prediction. Recent works investigate 3D gaze following [32, 30], gaze following for children [86], and object-aware gaze target prediction [89]. All these methods are fully-supervised. A concurrent work investigated semi-supervised gaze following using a saliency prediction model [66]. This approach does not leverage the strong priors from pretrained VL models, and its performance falls significantly behind our method.
@@
Semi-supervised Learning methods mostly adopt a teacher-student framework, which can be categorized into self-training [45, 10, 23, 82] and consistency regularization [44, 87, 19, 27]. Self-training methods generate pseudo annotations from teacher models trained with labeled data, while consistency regularization applies a consistency loss between the student and teacher model outputs, updating the teacher model gradually during training. Due to task differences, most semi-supervised recognition or segmentation methods are not directly applicable to gaze following [82, 65, 19, 37, 49]. However, some general methods without task-specific operations [44, 87] are applicable with appropriate modifications (see Supplementary). In our method, we adopt a self-training pipeline using the VQA prior and diffusion-based annotation prior, and show that we can get further improvements by using applicable consistency regularization methods (e.g. Mean Teacher [87]) to enhance the diffusion model training.
@@
Vision-Language Models have prospered after the development of large language models (LLMs) [90, 15, 13, 56, 55], and have succeeded in tasks such as visual grounding, visual question answering (VQA), and image captioning [92, 38, 79, 47, 53]. They also do well in low-shot generalization tasks by “prompting” [67, 60, 62], where task instructions are given to a pretrained model to generate outputs useful for other tasks. Grad-CAM [76] visualizations localize the linguistic input effect on the visual input [48, 98] thus achieving a degree of interpretability.
@@
Diffusion Models. Denoising diffusion probabilistic models (DDPM) [81, 28] give state-of-the-art results on image synthesis [16, 70, 21, 43, 22]. They are also used in inverse problems [84], image editing [58], and adversarial purification [64]. These works treat images with added noise as intermediate steps of the reverse process, showing that the noise magnitude affects the information retained from the input. Recent work adopts diffusion models in semi-supervised learning. [27] trains a diffusion model for 3D object detection, whereas [97] trains a diffusion model for image generation with images pseudo-labeled by existing semi-supervised models. In contrast, we use diffusion model as annotation prior, to refine the initial noisy annotations into high-quality pseudo annotations.
@@ 3.1 Overview
Figure 2: (a) Overall pipeline. We compute Grad-CAM heatmaps for unlabeled images and train the diffusion model with a small human-labeled set (or with unlabeled images using Mean Teacher). The diffusion model refines the Grad-CAM heatmaps into pseudo-annotations. Both the pseudo-annotations and the human-labeled set are used to train a gaze following model. (b) Grad-CAM heatmap generation. Given an image with an overlaid person bounding box, we “prompt" a pretrained VQA model with a gaze question and compute the Grad-CAM heatmap from the answer. (c) Grad-CAM refinement. We perturb the Grad-CAM heatmaps with Gaussian noise and pass through the reverse diffusion process to generate the refined pseudo-annotations.
@@ 3.2 Grad-CAM heatmap extraction
We use OFA [92], a transformer-based large pretrained VL model as the VQA model. Fig. 2b illustrates the procedure for Grad-CAM heatmap generation. Given an image $\bm{I}$ and the head bounding box of a person $\bm{l}_{h}$, we use Mask R-CNN [25] for person detection and find the bounding box that maximally overlaps with $\bm{l}_{h}$. We overlay this bounding box on top of the input image and give it as input to the VQA model along with a “prompt” in the form of a question: “What is the person in the green bounding box looking at?".
OFA has an encoder and a decoder. It generates a sequence of words as the answer. We compute Grad-CAM heatmaps $\mathcal{G}=\{\bm{g}_{j}\}_{j=1}^{N_{u}}$ on the decoder cross-attention weights between the input query and the image patch tokens based on [9] after selecting the noun from the answer. Details are in the supplementary.
@@ 3.3 Diffusion Model Training
Fig. 3(a) shows the training procedure. At each iteration, we sample a random time step $t$ from $U[1,T]$ to generate a noisy input $\mathbf{h}_{t}$ using Eq. 4. We feed $\mathbf{h}_{t}$ to the diffusion model for single-step denoising, and optimize with the loss of Eq. 5. We condition the diffusion model with the gaze feature $\mathbf{c}$ extracted from the input triplet $\bm{X}$ (input image, head crop, and head location mask) to output a gaze heatmap grounded on the gazing person; the parametrization of the U-Net denoiser becomes $\mathbf{h}_{0}^{\theta}(\mathbf{h}_{t},t,\mathbf{c})$. The feature extractor follows the structure of the VAT model [12], which takes the input triplet $\bm{X}$ and outputs the extracted features concatenated from two pathways for encoding the scene and gaze features. Details of the VAT structure are shown in Supplementary. Following [20], we add the extracted features as conditional features to each layer in the U-Net denoiser with the necessary transposed convolution layers to match the feature size of U-Net. The diffusion model can also be further improved by training with unlabeled data using Mean Teacher (MT) [87], a general-purpose semi-supervised learning method. As a consistency regularization method, MT enforces consistencies between the outputs from the teacher and student models fed with differently perturbed input. The teacher model has the same structure as the student model and is updated with an Exponential Moving Average (EMA) of the student model weights during training. In our case, we treat the diffusion model as the student model and apply different color jittering and head bounding box jittering to introduce different perturbations to the teacher and student model input. During training, the diffusion model is trained with the loss on labeled data and an additional consistency loss between the output heatmaps from the teacher and student diffusion models on all data. We demonstrate in our experiments that this further improves the gaze following results.
@@ 3.4 Heatmap Refinement using Diffusion Models
The refinement process is shown in Fig. 3(b). Given a noisy Grad-CAM heatmap $\bm{g}$ of an unsupervised input sample $\bm{X}$, we first add an appropriate amount of noise using the forward diffusion process $\hat{\mathbf{h}}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{g}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ according to Eq. 4. The added noise smoothes out the noise artifacts of the Grad-CAM heatmaps, while preserving the information of the highlighted regions. Then, by iteratively applying the denoising steps in Eq. 3, we get samples of the final heatmap $\hat{\mathbf{h}}_{0}$ that are geometrically and contextually similar to the human-labeled heatmap $\mathbf{h}_{0}$ while retaining gaze target information from the noisy Grad-CAM input. During the denoising process, we used DDIM [83] to sample the next step input using the predicted $\mathbf{h}_{0}$ from the U-Net.
@@ 4.1 Setting and Implementation Details
Implementation Details. We used a diffusion model with a linear noise schedule and $T{=}500$ steps. The heatmap size was $64{\times}64$. We used a batch size of 80 and a learning rate of $2.5\times 10^{-4}$ in the semi-supervised learning experiments. For training the diffusion model on labeled data, we used a learning rate of $5\times 10^{-5}$. When refining Grad-CAM heatmaps, we used DDIM [83] for inference.
@@ 4.2 Semi-supervised Training Results
We used VAT as the student model and compared our method of generating pseudo annotations with several alternatives [^1] Implementation details of the baselines are provided in the supplementary material.: 1) Semi-VAT: the pseudo-annotations for the unlabeled data were generated by the VAT model trained with labeled images. 2) Semi-VAT-GC: We created VAT-GC by adding Grad-CAM heatmaps as conditional inputs to the VAT model and modifying the VAT model accordingly. VAT-GC was trained on the labeled data to generate pseudo annotations for the unlabeled data. 3) VAT-MT: We use the Mean Teacher (MT) method [87] to train the VAT model. The first two baselines are self-training methods, while the 3rd belongs to the consistency regularization category.
@@ 4.3.2 Added Noise Magnitude Effects
We further visualized the outputs from the same diffusion model when adding noise at different time steps in Fig. 5. In the top 2 rows, the Grad-CAM heatmaps highlight correct locations, while in the bottom 2 rows, they are noisy or highlight incorrect locations. In all these cases, when adding noise at the 100th step, the outputs are located on the highlighted region of the Grad-CAM heatmap. When adding noise at the 400th step, the large magnitude of added noise leads to outputs similar to sampling directly from Gaussian noise without using the Grad-CAM heatmaps. Adding noise at step 250 appears to be the best trade-off.
@@ 4.3.3 The Effect of the Grad-CAM Heatmaps
In this section, we test how much guidance signal is provided by the VQA priors. This cannot be directly tested in the semi-supervised scenario, so we tested by using the Grad-CAM heatmaps as a direct input to a fully supervised gaze following model. We trained the teacher models of Sec. 4.2 using different amounts of labeled data only, and introduce a new baseline: Diffusion-GC, where the diffusion model trained on labeled images was used to refine the Grad-CAM heatmaps computed directly on the test set. Note that VAT-GC and Diffusion-GC require the Grad-CAM heatmaps of the test images during inference, which may not be feasible in an online setting as it necessitates access to a large vision-language model.
Tab. 4 presents the results. VAT-GC outperforms VAT in all cases, and Diffusion-GC performs best in the distance metrics. Both the VAT and diffusion model show large performance boosts when Grad-CAM heatmaps are directly used as input, which offers evidence that the VQA Grad-CAM heatmaps bring strong prior knowledge. As illustrated in Sec. 4.2, the lower AUC for Diffusion and Diffusion-GC can be attributed to the AUC on GazeFollow being evaluated with group-level annotations and favoring heatmaps spanning larger areas. We analyzed this in more detail in the supplementary material.